# Cosmic muon flux attenuation methods for superconducting qubit experiments

### v4

@@ 4.2 Results and data interpretation
In Fig. 8, we plot the energy spectrum up to 500 keV obtained during the simultaneous 16.28 hours run of the vertical and horizontal detectors. We observe a noticeable difference in the shape of both spectra, with a bump in counts in the horizontal detector between 140 keV and 360 keV, while the vertical detector has a flat count rate with only a peak at 200 keV due to the shining of the LED through the fiber optic.
@@
Muons with energies on the order of 1-100 GeV cross materials with a rather linear track and release energy at a constant rate. Hence, the energy spectra we are presenting are quite straightforward to interpret: in the vertical mounting, due to the angular dependence of the muon flux, a lower overall count rate is expected, since there is less available surface orthogonal to the flux. However, the horizontal position ensures that, statistically, most of the muons track will be short and in a confined energy range, hence the observed muon bump. The vertical geometry allows for longer tracks and more energy release for cosmic rays coming at orthogonal angles with respect to the ground.
@@
As a final remark, the mounting orientation of the chip can have a significant effect in mitigating the impact of ionizing radiation on quantum processors. However, it is unclear at present if it would be preferrable to have a lower count rate, but with more disruptive energy depositions, like in the vertical mounting scenario, or if it is better to minimize high energy releases, but endure more frequent particle interactions with the chip as in the horizontal mounting.
@@ 5 Conclusions
The first method consists in locating qubit experiments in a shallow underground location, which could be as low as 10 m below surface. This small depth already provides strong protection against neutrons, and it reduces the cosmic muon flux by a factor 2. Deeper sites at 100 m depth would observe a reduction of cosmic muon flux with factors on the order of 30. We have validated both methods using a home-built, portable cosmic muon counter consisting of two plastic scintillators integrated in a coincidence counting electronics circuit. The data collected by this portable detector has been cross-checked through Geant4 simulations.
The second method is linked with the mounting orientation of the chip in the dilution refrigerator. We employ two identical germanium wafers equipped with a particle sensor in a laboratory above ground to show how the vertical orientation of the chip reduces the count rate by 36 above 100 keV. However, this comes at the expense of enduring a higher count rate at large energy depositions within the substrate. In the present day, since the energy range that is most detrimental to qubit performances has not yet been determined, it is unclear which mounting orientation should be preferred.
@@ Acknowledgements
Data Availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Author Declarations
Conflict of interest
The authors have no conflicts to disclose.

### v6

@@ 4.2 Results and data interpretation
In Fig. 8, we plot the energy spectrum up to 500 keV obtained during the simultaneous 16.28 hours run of the vertical and horizontal detectors. We observe a noticeable difference in the shape of both spectra, with a bump in counts in the horizontal detector between 140 keV and 360 keV, while the vertical detector has a flat count rate with only a peak at 200 keV due to the shining of the LED through the optical fiber. We decided to show this LED peak, since these are events induced by particles and, as such, are not straightforward to remove from the dataset.
@@
Muons with energies on the order of 1-100 GeV cross materials along a rather linear track and release energy at a constant rate. Hence, the energy spectra we are presenting are quite straightforward to interpret: in the vertical mounting, due to the angular dependence of the muon flux, a lower overall count rate is expected, since there is less available surface orthogonal to the flux. However, the horizontal position ensures that, statistically, most of the muons track will be short and in a confined energy range, hence the observed muon bump. The vertical geometry allows for longer tracks and more energy release for cosmic rays coming at orthogonal angles with respect to the ground, or, equivalently, at shallow zenith angles.
@@
As a final remark, the mounting orientation of the chip can have a significant effect in mitigating the impact of ionizing radiation on quantum processors. However, it is unclear at present if it would be preferable to have a lower count rate, but with more disruptive energy depositions, like in the vertical mounting scenario, or if it is better to minimize high energy releases, but endure more frequent particle interactions with the chip as in the horizontal mounting.
@@ 5 Conclusions
The first method consists in locating qubit experiments in shallow underground sites, which could be as low as 10 m below surface. This small depth already provides strong protection against neutrons, and it reduces the cosmic muon flux by a factor 2. At deeper sites at 100 m depth we observe a reduction of cosmic muon flux with factors on the order of 30. Notably, the shallow sites identified in this work are located in urban areas with far greater accessibility than remote, deep underground laboratories, providing potential sites in which to install future, large-scale quantum processors. We have validated this method using a home-built, portable cosmic muon counter consisting of two plastic scintillators integrated in a coincidence counting electronics circuit. The data collected by this portable detector has been cross-checked through Geant4 simulations.
The second method is linked with the mounting orientation of the chip inside the dilution refrigerator. We employ two identical germanium wafers equipped with a particle sensor in a laboratory above ground to show how the vertical orientation of the chip reduces the count rate by 36 above 100 keV. However, this comes at the expense of enduring a higher count rate at large energy depositions within the substrate. In the present day, since the energy range that is most detrimental to qubit performances has not yet been determined, it is unclear which mounting orientation should be preferred.